Maria plans to invest a total of $7500 between two types of stocks, one yielding 6% interest and the other 15%. How much does she invest in each type of stock if the total interest gained annually is $585?
Answer by ankor@dixie-net.com(22740)

Maria plans to invest a total of $7500 between two types of stocks, one yielding 6% interest and the other 15%.
How much does she invest in each type of stock if the total interest gained annually is $585?
:
let x = amt invested at 15%
The total investment is $7500, therefore:
(7500-x) = amt invested at 6%
:
.06(7500-x) + .15x = 585
450 - .06x + .15x = 585
-.06x + .15x = 585 - 450
.09x = 135
x = 135/.09
x = $1500 invested at 15%
then
7500 - 1500 = $6000 invested at 6%
:
:
Confirm this by finding the actual interest from each investment
.06*6000 = $360
.15*1500 - $225
--------------------
total int: $585